# hard problem of consciousness

### v2

@@ Abstract
We tackle the hard problem of consciousness taking the naturally selected, self-organising, embodied organism as our starting point. We provide a mathematical formalism describing how biological systems self-organise to hierarchically interpret unlabelled sensory information according to valence and specific needs. Such interpretations imply behavioural policies which can only be differentiated from each other by the qualitative aspect of information processing. Selection pressures favour systems that can intervene in the world to achieve homeostatic and reproductive goals. Quality is a property arising in such systems to link cause to affect to motivate real world interventions. This produces a range of qualitative classifiers (interoceptive and exteroceptive) that motivate specific actions and determine priorities and preferences. Building upon the seminal distinction between access and phenomenal consciousness, our radical claim here is that phenomenal consciousness without access consciousness is likely very common, but the reverse is implausible. To put it provocatively: death grounds meaning, and Nature does not like zombies. We formally describe the multilayered architecture of self-organisation from rocks to Einstein, illustrating how our argument applies in the real world. We claim that access consciousness at the human level is impossible without the ability to hierarchically model i) the self, ii) the world/others and iii) the self as modelled by others. Phenomenal consciousness is therefore required for human-level functionality. Our proposal lays the foundations of a formal science of consciousness, deeply connected with natural selection rather than abstract thinking, closer to human fact than zombie fiction.
@@
hard problem of consciousness
@@ 1 Introduction
Why is anything conscious? Both biological and other physical [^1] Physical here just means non-biological. We are not suggesting biological systems are non-physical. systems process information, yet it seems that humans consciously experience in addition to merely processing information. Why? Living organisms are constantly processing self and world related information to survive in an ever-changing world. Human bodies share with all other physical systems the property of being instantiated in time and space, (e.g., our body occupies a given position and volume in space at a given time). Yet, unlike physical systems, biological, living systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation [nagel1974, chalmers1995].
Initially formally introduced in the field of cybernetics [ashby1947, foerster1960] the notion of self-organisation has been applied to various disciplines including physics [haken1983], biology [camazine2003, bell2007] and neuroscience [kelso1997, Friston2010, tognoli2014]. Self-organization is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment [camazine2001, seeley2002, rosas2018].
@@
This question is the subject of long-standing debates across disciplines [seth2022]. One highly influential view is that consciousness has two aspects [seth2022]. The first is functional, by which we mean the ability to access and communicate information [block1995]. How information processing is linked to consciousness is the “easy problem” of consciousness [chalmers1995]. The second aspect is “what it is like” to consciously experience information processing, or phenomenal consciousness [chalmers1995, block1995, nagel1974, gallagher2021, fuchs2017]. This doesn’t mean just global states like being awake, but more specific local states like smelling a cup of coffee. These local contents or “qualia” are characterised by what it is like to be in them [seth2022]. It is unclear the extent to which functional and phenomenal aspects are independent. David Chalmers has influentially suggested that it may be possible to construct a “philosophical zombie” which acts in every way like a person but has no qualia [block1995]. For example, a thermostat certainly detects heat and so processes information, but there presumably is not anything it is like to be a thermostat. Hence the question “why is anything conscious” may be understood as “why is there sometimes a qualitative aspect to information processing?”. This is the “hard problem” of consciousness.
The hard problem has sparked a substantial body of work, detailed discussion of which [seth2022, northoff2014] lies beyond the scope of our paper. Rather, in what follows we build upon the useful distinction between lower and higher order theories of information processing in relation to conscious experiences.
For example, higher order thought theory (HOT) [rosenthal2005, brown2019] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states. Lower order states may include emotions and perceptions, while higher order meta-representations reflect upon those. The link between the two may explain something of the phenomenal character of states. Sense data is processed by the body resulting in lower order mental states, and then meta representations of those is where we might find more abstract conceptual or thought-like contents of consciousness.
The distinction between lower and higher orders is a good starting point to understand why most biological information processing goes on “in the dark”. Yet, the interesting question in our view is why do these lower order states arise? Can they occur in the absence of subjective, qualitative experience? It is widely agreed that higher order mental states such as desire to drink a coffee is accompanied by a qualitative aspect. But is there something it is like to “be in the dark”? That is: to process information at the lower bodily levels?
@@
If an explanation of how consciousness functions is to reveal why information processing sometimes has a qualitative aspect, then it must explain how we get local states [thompson2016], not begin by assuming local states. We need to go a level down and begin at the level of the embodied organism. Longtime considered a fringe approach, the embodied cognition paradigm [varela2016] has recently gained substantial influence in cognitive science and philosophy [ciaunica2023, levin2024, seth2018]. The idea is that instead of considering the body as a mere device designed to fuel and contain the mind (a device that can be replaced with a vat or a robot, for example), one must consider the mind as serving the self-sustaining needs of a surviving body.
@@
One promising approach to this question is to try to rigorously define consciousness from first principles and show that some aspects of functional consciousness depend on phenomenal consciousness in a manner that makes zombies impossible, “dissolving” the hard problem by showing the phenomenal to be intrinsically functional [^2] Note that we are far from the first to claim that the phenomenal is functional. For example, the proposed Conscious Turing Machine [blum2020] based on Global Workspace Theory [baars1997], and constructivist approaches to artificial intelligence [wang2020] take similar position. It is our explanation of how and why that is novel.. Unifying them. We would need to establish axioms that hold in every conceivable environment and show that it is impossible to have the function of consciousness without the subjective experience of it. Such an answer must explain how consciousness functions, and why most information processing goes on “in the dark”.
This paper lays the grounds of a computational model formalizing the link between lower and higher levels of information processing in relation to conscious experiences in a way that is compatible with the basic principles of embodied cognition and enactivism. Enactivism is roughly the view that information processing arises through a dynamic interaction between an acting organism and its environment [thompson2007].
@@
Now, the notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression [piccinini2015, piccinini2021]. Here we define computation not in terms of abstract symbol shuffling and representation, but in concrete mechanistic terms where ‘information processing’ refers to the actual physical transition of a system from one state to another; a system such as a human body.
Pancomputationalism is the idea that all dynamic systems are constituted by computation, which for our purposes is defined as above. This paper builds upon the formalism of pancomputational enactivism [bennett2024a]. To answer “why is anything conscious” we proceed in two steps. First, we develop a formalism in which lower and higher order theories of consciousness, and phenomenal and access consciousness, are all derived from first principles. Each follows because of scaling natural selection pressures and the ability to adapt [^3] This is loosely inspired by a scale-based framing of machine learning [sutton2019] - we are applying Sutton’s “bitter lesson” to biology.. Second, we provide an argument building on that to dissolve the hard problem of consciousness, answering why there must a qualitative aspect to information processing and in what circumstances.
@@
This paper is organised as follows. It should be noted first that the appendix at the end of this paper is a list of formal mathematical definitions, which the reader may refer to as needed. For readers without formal background, the appendix can be skipped. In Section 2 we extend separately published arguments, experiments and proof results [bennett2024a, bennett2023b, bennett2023c], developing a model of self-organising systems that holds across all conceivable environments. In Sections 3 and 4 we explain how this model formalises relevance realisation and unifies lower and higher order theories. Rather than assuming abstract objects are primary and trying to learn causal relations between them, this assumes valence is primary and that abstract objects are constructed to classify cause and anticipate valence. This takes us from attraction and repulsion to contentless states, through lower order thoughts to higher order meta representations. We call this the psychophysical principle of causality. In Section 5 we extend previous work [bennett2023c] on causal learning and the development of first (1ST), second (2ND) and third (3RD) order selves, and explain how they are constructed as a direct consequence of scaling the ability of a self-organising system to adapt with natural selection pressures. In Sections 6 and 7 we explain how subjective experience requires a 1ST order self, and conversely why a 1ST order self implies there is something it is like to be an organism that has a 1ST order self. We then argue that access consciousness requires the ability to communicate meaning, which requires both 1ST and 2ND order selves, and show how a philosophical zombie is impossible. In Section 8 we describe the development of consciousness as we scale up the capacity to adapt with natural selection pressures. We discuss examples of each stage from nematodes to humans. Section 9 is concluding remarks.
@@ 2 Back to Foundations
Rather than presupposing local states or an abstraction layer, we start from first principles to formalise all conceivable environments [bennett2024a, bennett2024b], as in definition 1. This is based on two assumptions:
@@
Axiom 2: Where things differ, we have different states of the environment.
Hence, the environment is a set $\Phi$ of contentless global states $\Phi$. We define a declarative “program” as a set of states and treat the states themselves as irreducible [^4] We do not concern ourselves with imperative programs, because there is an equivalence between declarative and imperative [howard1980].. The set $P$ contains all such programs. The powerset of $P$ is every aspect of the environment, because every aspect of any conceivable environment must be set of such programs, which is a subset of $P$.
Programs return true or false, so a program is a “fact” if it is true, and the current state of the environment is the set of all “facts”. Truth is determined with respect to states, and states are assumed to be mutually exclusive. For example, if time is one way in which things differ, then there is one state at a time. It follows that every aspect of the environment is a set of programs, and those programs are facts relative to any physical state they contain. An aspect of the environment is realized by a state if the programs it contains are facts.
When we say a state is contentless, we mean we assume nothing about what states contain or are. Their only significance is in how they relate to one another, and each such relation is a program. This is intentionally like to post-structuralism [eacho2020, hayles2005], in which the only meaning words have is given via their relations to each other. This is why previous work claims it applies to all conceivable environments [bennett2024a].
Universality Claim: Axioms 1 and 2 hold for every conceivable environment.
Intuitively, one might understand the set of all programs (the powerset of all states) as the set of all relations between these otherwise meaningless states.
@@ 2.1 Natural Selection and Embodiment
We then assume a process of natural selection. This produces embodied organisms.
@@
A set of facts can represent anything [^5] Some may object, pointing out that this ignores composition. However, the application of a function is a fact regardless of whether it takes another function as input., hence the body of an organism must be a set of programs. We call this set a vocabulary, and it is formalised in definition 2 (see Appendix).
@@
The vocabulary can be seen as hardware and software operating together. Piccinini [piccinini2015, piccinini2021] identifies two sorts of computation: abstract and concrete. Abstract computations mean or represent whatever we say they do. In contrast concrete computations are physical. They are the action of physical systems over time. The programs we talk about here are concrete, physical computations in that they are the relations between physical states. It follows that the computations a body performs are also concrete, and so no two organisms have the same vocabulary, because that would mean they are the same body.
@@
This might seem inconvenient, as intuitively we might prefer to think about bodies as an abstract ‘type’. However, doing so would undermine any claim we might make, as earlier results in artificial intelligence (AI) show [bennett2023c, bennett2024c]. The meaning of abstractions depends upon interpretation, much like the behavior of software is determined by the abstraction layer that “interprets” it. At its least abstract software is machine code [^6] A “word” of machine code triggers a mechanistic process hardwired in the CPU by the human design. For example a word may copy the 32 bit value stored at memory address $X$ into the 32 bit register $Y$, which the next line of machine code “adds” to register $Z$ by looping over each bit in an “adder” circuit.. Machine code is interpreted by hardware, and that hardware determines every aspect of what that machine code does. A word of machine code is a mechanical trigger that only “means” whatever we have designed the hardware to do when we input that word. That hardware is an abstraction layer in which the software exists (including “higher level” abstraction layers like the Python interpreter). In other words, what we call software is nothing more than the state of hardware [bennett2024a]. This has undermined all but the most subjective of claims regarding the behavior of theorised, software superintelligence [hutter2010, leike2015, leike2018]. It is a flaw in the very idea of intelligent software. The distinction between software mind and hardware embodiment is subsequently called computational dualism [bennett2024a], because it is reminiscent of how traditional cartesianism conceived of a mental substance distinct from physical.
@@
Here we suggest that must avoid computational dualism. If we are to dissolve the hard problem by starting with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states, then we cannot presuppose organisms use particular abstractions. Instead we must explain why particular abstractions are formed, in terms of contentless states, as the product of interaction between an organism and its environment [thompson2007, varela2016, rolla2021]. This process is sometimes called “relevance realisation” [vervaeke2012, vervaeke2013a, vervaeke2013b].
@@
This means that a body can occupy only a finite number of states, which is formalised in definition 2. A body is a vocabulary $\mathfrak{v}\subset P$ with finitely many elements. $\mathfrak{v}$ implies a formal language $L_{\mathfrak{v}}$ of interaction between body and environment. A statement in this formal language is an aspect of the environment. A set of “programs” from our vocabulary. Statements have truth conditions with respect to environmental states, and every statement has an extension. The extension of a statement is the set of all statements which are supersets of the first statement (the set of all other statements by which the first statement is implied). Extension is important we can relate statements by their truth conditions, forming a lattice. Most importantly, we avoid the distinction between software and hardware and so avoid computational dualism.

### v3

# xxxxx, xxxxx, xxxx

@@ Abstract
We tackle the hard problem of consciousness taking the naturally selected, embodied organism as our starting point. We provide a formalism describing how biological systems self-organise to hierarchically interpret unlabelled sensory information according to valence. Such interpretations imply behavioural policies which are differentiated from each other only by the qualitative aspect of information processing. Natural selection favours systems that intervene in the world to achieve homeostatic and reproductive goals. Quality is a property arising in such systems to link cause to affect to motivate interventions. This produces interoceptive and exteroceptive classifiers and determines priorities. In formalising the seminal distinction between access and phenomenal consciousness, we claim that access consciousness at the human level requires the ability to hierarchically model i) the self, ii) the world/others and iii) the self as modelled by others, and that this requires phenomenal consciousness. Phenomenal without access consciousness is likely common, but the reverse is implausible. To put it provocatively: death grounds meaning, and Nature does not like zombies. We then describe the multilayered architecture of self-organisation from rocks to Einstein, illustrating how our argument applies. Our proposal lays the foundation of a formal science of consciousness, closer to human fact than zombie fiction.
@@
hard problem, consciousness
journal: rsif
{fmtext}
@@ 1 Introduction
Why is anything conscious? Both biological and other physical [^1] Physical here just means non-biological. We are not suggesting biological systems are non-physical. systems process information, yet humans consciously experience as well as process information. Why? Living organisms are constantly processing self and world related information to survive in an ever-changing world. Human bodies share with all other physical systems the property of being instantiated in time and space (e.g. occupy space). Yet, unlike physical systems, biological systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation [nagel1974, chalmers1995].
Initially formally introduced in the field of cybernetics [ashby1947, foerster1960] the notion of self-organisation has been applied to various disciplines including physics [haken1983], biology [camazine2003, bell2007] and neuroscience [kelso1997, Friston2010, tognoli2014]. Self-organisation is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment [camazine2001, seeley2002, rosas2018].
@@
This question is the subject of long-standing debate [seth2022]. One highly influential view is that consciousness has two aspects [seth2022]. The first is functional, by which we mean the ability to access and communicate information [block1995]. How function relates to consciousness is considered the “easy problem” of consciousness [chalmers1995]. The second aspect is “what it is like” to consciously experience information processing, or phenomenal consciousness [chalmers1995, block1995, nagel1974, gallagher2021, fuchs2017]. This doesn’t mean just global states like being awake, but more local states like smelling a cup of coffee. These local contents or “qualia” are characterised by what it is like to be in them [seth2022]. It is unclear the extent to which functional and phenomenal aspects are independent. David Chalmers has influentially suggested that it may be possible to construct a “philosophical zombie” which acts in every way like a person but has no qualia [block1995]. For example, a thermostat certainly detects heat and so processes information, but there presumably is not anything it is like to be a thermostat. Hence the question “why is anything conscious” may be understood as “why is there sometimes a qualitative aspect to information processing?”. This is the “hard problem” of consciousness.
The hard problem has sparked a substantial body of work, detailed discussion of which [seth2022, northoff2014] lies beyond the scope of our paper. In what follows we build upon the distinction between lower and higher order theories of information processing in relation to conscious experiences.
Higher order thought theory (HOT) [rosenthal2005, brown2019] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states. Lower order states may include emotions and perceptions, while higher order meta-representations reflect upon those. The link between the two may explain something of the phenomenal character of states. Sense data is processed by the body resulting in lower order mental states, and then meta representations of those is where we might find more abstract conceptual or thought-like contents of consciousness.
HOTs are a good starting point to understand why most biological information processing goes on “in the dark”, but why do these lower order states arise? Can they occur in the absence of subjective, qualitative experience? Is there something it is like to “be in the dark” processing information at the lower bodily levels?
@@
Longtime considered a fringe approach, the embodied cognition paradigm [varela2016] has recently gained substantial influence in cognitive science and philosophy [ciaunica2023, levin2024, seth2018]. The idea is that instead of considering the body as a mere device designed to fuel and contain the mind (a device that can be replaced with a vat or a robot, for example), one must consider the mind as serving the self-sustaining needs of a surviving body.
@@
Our approach is to try to rigorously define consciousness from first principles and show that some aspects of functional consciousness depends on phenomenal consciousness in a manner that makes zombies impossible, “dissolving” the hard problem by showing the phenomenal to be intrinsically functional [^2] Note that we are far from the first to claim that the phenomenal is functional. For example, the proposed Conscious Turing Machine [blum2020] based on Global Workspace Theory [baars1997], and constructivist approaches to artificial intelligence [wang2020] take similar position. It is our explanation of how and why that is novel.. We establish axioms that hold in every conceivable environment and show that it is impossible to have the function of consciousness without the subjective experience of it, and why most information processing goes on “in the dark”.
@@
Enactivism is roughly the view that information processing arises through interaction between an organism and its environment [thompson2007], and it is considered by some to be incompatible with narrow definitions of computation. The notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression [piccinini2015, piccinini2021]. Here we define computation not in terms of abstract symbol shuffling and representation, but in concrete mechanistic terms where ‘information processing’ refers to the actual physical transition of a system from one state to another, so it is compatible with enactivism. Pancomputationalism is the idea that all systems are computational. We build upon a formalism called pancomputational enactivism [bennett2024a].
@@
To answer “why is anything conscious”, we use this formalism to show how lower and higher order theories, phenomenal and access consciousness all follow from first principles, scaling natural selection pressures and the ability to adapt [^3] This is loosely inspired by a scale-based framing of machine learning [sutton2019] - we are applying Sutton’s “bitter lesson” to biology.. First, in Section 2 we integrate and extend separately published works of narrower scope [bennett2024a, bennett2023b, bennett2023c] to justify our model of self-organising systems. In Sections 3 and 4 we explain how this model formalises relevance realisation and unifies lower and higher order theories. Rather than assuming abstract objects are primary and trying to learn causal relations between them, this assumes valence is primary and that abstract objects are constructed to classify cause and anticipate valence. This takes us from attraction and repulsion to physical states, through lower order thoughts to higher order meta representations. We call this the psychophysical principle of causality. In Section 5 we extend previous work [bennett2023c] on causal learning and the development of first (1ST), second (2ND) and third (3RD) order selves, and explain how they are a consequence of scaling the ability of a self-organising system to adapt with natural selection pressures. In Sections 6 and 7 we explain how subjective experience requires a 1ST order self, and conversely why a 1ST order self implies there is something it is like to be an organism that has a 1ST order self. We then argue that access consciousness requires the ability to communicate meaning, which requires both 1ST and 2ND order selves, and show how a philosophical zombie is impossible. In Section 8 we describe the development of consciousness as we scale up the capacity to adapt with natural selection pressures, with examples from nematodes to humans.
@@ 2 Back to Foundations
To meet length restrictions, the full mathematical definitions and proofs have been placed in the appendices [bennett2024appendices]. The core arguments should be understandable without the math, but for convenience here is a quick reference guide:
  1. 1.
    A set $\Phi$ is assumed, whose elements are called states. A declarative program is set $f$ of states, and $f$ is true relative to $\phi\in\Phi$ iff $\phi\in f$. $P$ is the set of all programs. Embodiment is formalised as an abstraction layer, for which we single out finite set $\mathfrak{v}\in 2^{\Phi}$ called a vocabulary. A choice of vocabulary $\mathfrak{v}$ implies a formal language $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}$.
  2. 2.
    Members of $L_{\mathfrak{v}}$ are called statements, and a statement $l$ is considered true relative to a state $\phi$ iff $\phi\in\bigcup l$. A completion of a statement $l_{1}$ is any statement $l_{2}$ s.t. $l_{1}\subseteq l_{2}$. The extension of a statement $l$ is the set $E_{l}$ of all its completions, s.t. $x\in{L_{\mathfrak{v}}}$ is $E_{x}=\{y\in{L_{\mathfrak{v}}}:x\subseteq y\}$. The extension of a set $X\subseteq{L_{\mathfrak{v}}}$ of statements is $E_{X}=\bigcup\limits_{x\in X}E_{x}$.
  3. 3.
    A $\mathfrak{v}$-task $\alpha$ is a pair $\langle I_{\alpha},O_{\alpha}\rangle$ where $I_{\alpha}\subset L_{\mathfrak{v}}$ are the inputs, $E_{I_{\alpha}}$ is the outputs available given ${I_{\alpha}}$, and $O_{\alpha}\subset E_{I_{\alpha}}$ are the correct outputs. All the tasks an abstraction layer tasks form a hierarchy. If $\alpha\sqsubset\omega$ then $\omega$ is a parent of $\alpha$, meaning $I_{\alpha}\subset I_{\omega}$ and $O_{\alpha}\subseteq O_{\omega}$. A parent is higher level than its children.
  4. 4.
    A policy is a statement $\pi$ and a policy is correct for $\alpha$ iff $\pi\in{\Pi_{\alpha}}=\{\pi\in L_{\mathfrak{v}}:{E}_{I_{\alpha}}\cap E_{\pi}={O_{\alpha}}\}$. Given input $i\in I_{\alpha}$ and a policy $\pi$, to infer is to choose $o\in E_{i}$. A task is complete if $o\in O_{\alpha}$. Given a task definition $\alpha$ s.t. $\alpha\sqsubset\omega$, to learn is to choose a policy $\pi\in\Pi_{\alpha}$, and $\omega$ has been learned from $\alpha$ if one chooses $\pi\in\Pi_{\omega}$.
  5. 5.
    A causal identity for $\mathfrak{b}$ is a policy classifying the causal interventions undertaken by $\mathfrak{b}$. Subscript denotes the system that learned the causal identity, while superscript denotes the object it classifies. For example, $c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes a causal identity learned by $\mathfrak{a}$ to classify $\mathfrak{b}$’s interventions, and $c^{\mathfrak{ba}}_{\mathfrak{a}}\subset c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$.
We begin by formalising all conceivable environments [bennett2024a, bennett2024b] from first principles:
@@
Axiom 2: Where things change or differ, we have different states of the environment.
Hence, the environment is a set $\Phi$ of mutually exclusive states. We call a set of states a declarative “program” [^4] We do not concern ourselves with imperative programs, because there is an equivalence between declarative and imperative [howard1980].. A program is “true” relative to the states it contains. For example, if time is one dimension along which things change, then there is one state at a time. A “fact” is a program which contains the state at the present time. It follows that every aspect of the environment is a set of programs. An aspect of the environment is realised by (exists in) a state if the programs it contains are facts. Previous work argues this formalises every conceivable environment [bennett2024b].
To give some intuition as to why, note that we assume nothing of what states might be or contain. We assume no differences within a state, only between states, and those differences are the programs of which aspects are formed. After all, one can only interact with aspects of one’s environment. If one were to try and pinpoint what an aspect was made of, the answer would be deferred to other aspects. Any description that sought to fully encapsulate the semantics, truth or unmediated pure experience of a state would always be delayed or incomplete. This does not render such attempts vacuous, but speaks to their inherent contingency and conditionality, which is what this formalism attempts to capture.
@@ 2.1 Natural Selection and Embodiment
We assume a naturally selected, embodied organism. Every aspect of the environment is a set of facts [^5] Some may object, pointing out that this ignores composition. However, the application of a function is a fact regardless of whether it takes another function as input., hence the body of an organism that persists across multiple states must be a set of programs. We call this set a vocabulary, and it is formalised in definition 2 (see appendix [bennett2024appendices]).
@@
Piccinini [piccinini2015, piccinini2021] divides computation into abstract and concrete sorts. Abstract computations mean or represent whatever we say they do. In contrast concrete computations are the actions of physical systems. The declarative programs we describe are concrete. Each captures a single point of difference between physical states. As such no two organisms have the same vocabulary because that would mean they are the same body.
@@
This might seem inconvenient, as intuitively we might prefer to think about bodies as an abstract ‘type’. However, doing so would undermine any claim we might make, as earlier results in artificial intelligence (AI) show [bennett2023c, bennett2024c]. The meaning of abstractions depends upon interpretation, much like the behaviour of software is determined by the abstraction layer that “interprets” it. At its least abstract software is machine code [^6] A “word” of machine code triggers a mechanistic process hardwired in the CPU by the human design. For example a word may copy the 32 bit value stored at memory address $X$ into the 32 bit register $Y$, which the next line of machine code “adds” to register $Z$ by looping over each bit in an “adder” circuit.. Machine code is interpreted by hardware which determines every aspect of what that machine code does. A word of machine code is a mechanical trigger that only “means” whatever we have designed the hardware to do when we input that word. That hardware is an abstraction layer in which the software exists (including “higher level” abstraction layers like the Python interpreter). In other words, what we call software is nothing more than the state of hardware [bennett2024a]. This has undermined all but the most subjective of claims regarding the behaviour of theorised, software superintelligence [hutter2010, leike2015, leike2018]. It is a flaw in the very idea of intelligent software. The distinction between software mind and hardware embodiment is subsequently called computational dualism [bennett2024a], because it is reminiscent of how traditional Cartesianism conceived of a mental substance distinct from physical.
@@
If we are to dissolve the hard problem by starting with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states, then we cannot presuppose organisms use particular abstractions. Instead we must explain why particular abstractions are formed, in terms of contentless states, as the product of interaction between an organism and its environment [thompson2007, varela2016, rolla2021]. This process is sometimes called “relevance realisation” [vervaeke2012, vervaeke2013a, vervaeke2013b].
@@
This means that a body is realised by only a finite number of states (see def. 2). A body is a vocabulary $\mathfrak{v}\subset P$ with finitely many elements. $\mathfrak{v}$ implies a formal language $L_{\mathfrak{v}}$ of interaction between body and environment. A statement in this language is an aspect of the environment. A set of “programs” from our vocabulary. Statements have truth conditions with respect to environmental states, and every statement has an extension. The extension of a statement is the set of all statements which are supersets of the first statement (the set of all other statements by which the first statement is implied). Extension is important we can relate statements by their truth conditions, forming a lattice. Most importantly, we avoid the distinction between software and hardware and so avoid computational dualism.